OpenStudy (chosenmatt):

what is the sum? 10/19+3/19

OpenStudy (catlover5925):

10/19 is 0.526 and 3/19 is 0.157 so add that and you get 0.683

OpenStudy (chosenmatt):

i need help with finding answer in a fraction not decimal

OpenStudy (chosenmatt):

@iGreen

OpenStudy (igreen):

10 + 3 = ? @chosenmatt

OpenStudy (catlover5925):

ok so it would be 683/1000...is that a choice? im not sure if that is simplified

OpenStudy (chosenmatt):

13

OpenStudy (igreen):

Yep, so that gives us 13/19

OpenStudy (chosenmatt):

yay i had the right answer!!

OpenStudy (catlover5925):

awesome @iGreen

OpenStudy (igreen):

\(\dfrac{10}{19} + \dfrac{3}{19} \rightarrow \dfrac{10+3}{19} \rightarrow \dfrac{13}{19}\)